### v1

# Random Latent Exploration for Deep Reinforcement Learning

Abstract
The ability to efficiently explore high-dimensional state spaces is essential for the practical success of deep Reinforcement Learning (RL). This paper introduces a new exploration technique called Random Latent Exploration (RLE), that combines the strengths of bonus-based and noise-based (two popular approaches for effective exploration in deep RL) exploration strategies. RLE leverages the idea of perturbing rewards by adding structured random rewards to the original task rewards in certain (random) states of the environment, to encourage the agent to explore the environment during training. RLE is straightforward to implement and performs well in practice. To demonstrate the practical effectiveness of RLE, we evaluate it on the challenging Atari and IsaacGym benchmarks and show that RLE exhibits higher overall scores across all the tasks than other approaches.
Machine Learning, ICML
11footnotetext: Project website: https://srinathm1359.github.io/random-latent-exploration
1 Introduction
Reinforcement learning (RL) (Sutton & Barto, 2018) trains agents to maximize rewards through interactions with the environment. Since rewards can be delayed, focusing only on immediate rewards often leads to sub-optimal long-horizon strategies. Often, agents must sacrifice short-term rewards to discover higher rewards. Identifying actions that eventually result in higher rewards, known as the exploration problem, is a core challenge in RL.
Exploration is challenging because the current action’s effect may often be revealed only after many interactions with the environment. Exploration is well-studied(Amin et al., 2021) and the major approaches can be broadly categorized into two types: (i) Noise-based exploration (e.g., $\epsilon$-greedy, Boltzmann sampling, posterior sampling (Osband et al., 2016a, 2019; Fortunato et al., 2017; Ishfaq et al., 2021, 2023; Plappert et al., 2017)) and (ii) Bonus-based exploration (Bellemare et al., 2016; Pathak et al., 2017; Burda et al., 2018; Oudeyer & Kaplan, 2009; Pathak et al., 2019; Hong et al., 2018). Unfortunately, neither family of approaches consistently outperforms the other when performance is measured across a range of tasks with either discrete (Chen et al., 2022) or continuous (Schwarke et al., 2023) action spaces. Unsurprisingly, the choice of exploration strategy for a new task is intimately tied to the task characteristics that are difficult to determine in advance. Therefore, the current common practice is finding the best exploration scheme using the trial-and-error process of trying different strategies.
Noise-based exploration typically involves perturbing the policy’s parameters (such as the weights of the policy network (Fortunato et al., 2017; Plappert et al., 2017)), or the action output of the policy (e.g. $\varepsilon$-greedy). The added noise prevents the agent from generating the same trajectory repeatedly, thereby encouraging the exploration of different trajectories. Noise-based exploration is the go-to exploration scheme in deep RL due to its simplicity of implementation. However, such strategies are less effective in tasks requiring deep exploration than bonus-based exploration strategies (Osband et al., 2016a). One possible reason revealed by our experiments (see Section 4.1) is that commonly used perturbation strategies only affect the policy locally and, therefore, do not explore states far from the initial states.
Bonus-based strategies augment the task reward with an incentive (or a bonus) (e.g., prediction error (Pathak et al., 2017), visitation count (Bellemare et al., 2016), information gain (Houthooft et al., 2016)) that encourages the agent to explore far away states and thereby achieve deep exploration. Unfortunately, computing the bonus requires training an additional deep neural network. Furthermore, while bonus-based exploration outperforms noise-based exploration in a few hard-exploration tasks, when average performance is measured across a range of tasks (i.e., both easy and hard exploration problems), both strategies perform similarly (Taïga et al., 2019; Chen et al., 2022).
Instead, we hypothesize that it is easier to explore by training the agent to achieve a variety of diverse goals, i.e. exploring the space of goals. Following this intuition, in this work, we introduce a different exploration strategy where instead of injecting noise or adding bonuses to encourage exploration, we train the agent to achieve various goals from a goal space $\mathcal{Z}$. Our key intuition is that if the goals are designed such that different goals incentivize the agent to explore diverse parts of the environment, then training an agent to achieve these goals will lead it to explore a variety of states, eventually leading it to find and exploit high task rewards. Unfortunately, representing and selecting "the right goals" is challenging in practice, as they may be dependent on the underlying task in complex ways that the learner cannot anticipate in advance.
To circumvent the issue of finding the right goals to target, we propose Random Latent Exploration (RLE), where the agent’s policy is conditioned on random vectors sampled from a fixed distribution as goals, which define additional reward bonuses differing at each state based on the chosen vectors. In particular, the random vectors act as latent goals, each creating different reward functions that encourage the agent to reach different parts of the environment. By sampling enough random vectors during training, the agent is trained to pursue many different goals, thus resulting in deep exploration. Our experiments show that RLE leads to significantly more diverse and deeper trajectories than traditional noise-based methods, without the need for complex bonus computation. This makes RLE both easy to implement and effective in practice.
To show the effectiveness of our approach, we evaluated RLE in Atari—a popular discrete action space deep RL benchmark (Bellemare et al., 2013), and IsaacGym—a popular continuous control deep RL benchmark (Makoviychuk et al., 2021), each consisting of many different tasks with varying degrees of exploration difficulty. We implement our method on top of the popular RL algorithm, Proximal Policy Optimization (PPO) (Schulman et al., 2017) and compare it with PPO in other exploration strategies. Our experimental results demonstrate that RLE improves over standard PPO in Atari and IsaacGym. Furthermore, RLE also exhibits a higher aggregated score across all tasks in Atari than other exploration methods, including RND (Burda et al., 2019) and randomized value function strategies (Fortunato et al., 2017). Importantly, these improvements were obtained by simply adding RLE on the top of the base PPO implementation while only changing the learning rate (and discount rate for Atari). We use the same hyperparameters across all 57 Atari games and the same hyperparameters across all 9 IsaacGym tasks, thus highlighting the generality of our approach as a plug-in utility.
2 Preliminaries
Reinforcement Learning (RL).
RL is a popular paradigm for solving sequential decision-making problems (Sutton & Barto, 2018) where an agent operates in an unknown environment (Sutton & Barto, 2018) and aims to improve its performance through repeated interactions with the environment. At each round of interaction, the agent starts from an initial state $s_{0}$ of the environment and collects a trajectory. At each timestep $t$ within that trajectory, the agent perceives the state $s_{t}$, takes an action $a_{t}\sim\pi(.|s_{t})$ with its policy $\pi$, receives a task reward $r_{t}=r(s_{t},a_{t})$, and transitions to a next state $s_{t+1}$ until reaching terminal states, after which a new trajectory is initialized from $s_{0}$ and the above repeats. The goal of the agent is to learn a policy $\pi$ that maximizes expected return $\mathbb{E}_{\pi}\Big{[}\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})\Big{]}$ in a trajectory. A straightforward approach is to estimate the expected return of a policy by rolling out trajectories $(s_{0},a_{0},s_{1},\cdots,)$ through Monte Carlo sampling (Konda & Tsitsiklis, 1999), and then optimizing this to find the optimal policy, but unfortunately, the corresponding estimates are of high variance and thus often require a huge number of data. Thus, in practice, various RL algorithms learn a value function (or value network) $V^{\pi}$ from the interaction that approximates
$V^{\pi}(s_{0})\approx\mathbb{E}_{\pi}\Big{[}\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})\Big{]},$ (1)
and train the policy $\pi$ to maximize the value $V^{\pi}(s_{0})$ (e.g. using policy gradient).
Exploration. As the reward may be delayed and not immediately presented to the agent, the agent may need to take many actions and visit a sequence of states without rewards before it receives any learning signal (reward). As such, taking greedy action $a_{t}$ at each step that maximizes immediate reward $r(s_{t},a_{t})$ does not necessarily lead to a high return. Thus, RL algorithms require “exploring” states and actions that may lead to low immediate rewards but could potentially end up with high return in the long run. We refer to this process as exploration throughout this paper.
3 Our Method: Random Latent Exploration
Problem statement. Our goal in this paper is to create an exploration strategy that improves the standard action noise exploration method overall on the majority of tasks and domains (both discrete and continuous control) and yet is easy to implement. Toward that end, we provide a simple exploration method that combines the best of both worlds and bridges the approach of using reward bonuses (Bellemare et al., 2016; Burda et al., 2019) with noises or posterior sampling for exploration (Osband et al., 2016a, b; Fortunato et al., 2017).
Challenge: Ensuring diversity. The key challenge in noise-based exploration methods lies in their ability to generate diverse behaviors within the environment. Specifically, if noise-perturbed policies bear too much similarity, they will yield nearly identical trajectories $(s_{0},a_{0},r_{0},s_{1},\cdots)$, consequently limiting exploration to a narrow region of the environment. The primary technical challenge, then, is devising a method that ensures such diversity in the generated trajectories.
Approach: Randomized rewards. To overcome the above challenge, we propose enhancing trajectory diversity by altering the agent’s rewards, inspired by skill discovery methods (Eysenbach et al., 2018) that use varied reward functions to develop diverse skills resulting in different state visitation distribution. In particular, in every episode of training, we randomly perturb the given task reward function by adding a randomized reward to certain (random) parts of the state space. Then, we train policies to maximize the composite of randomized and task rewards. The key idea is that in every round of interaction, the random rewards added to certain random parts of the state space will incentivize the agent to visit those areas, and if these random areas are diverse enough, we will get diverse behaviors in the environment during training—thus incentivizing exploration. However, since the random rewards are repeatedly resampled and thus keep on changing during training to ensure stable and effective learning, both policies and value functions must be aware of the specific random reward function in use; otherwise, the changing reward functions will comprise a partially observable MDP (Kaelbling et al., 1998). We take inspiration from the Universal Value Function Approximator (UVFA) (Schaul et al., 2015), which trains networks based on varying goals. We adopt their approach by equating their goals to different reward functions, thus making the policy $\pi$ and the value function $V^{\pi}$ condition on the sampled reward function. This ensures that the random rewards no longer appear as noises to the policy. The remaining questions are:
  • How to implement the randomized reward functions?
  • How to make the policy and the value function condition on the sampled reward functions?
We outline our implementation of the above idea in the next section and defer full implementation details to Appendix B.
3.1 Algorithmic Implementation
Randomized reward functions. An effective approach to implementing randomized reward functions necessitates adherence to two fundamental principles. First, the randomized rewards require correlating with states; otherwise, they would appear as white noises, which does not help exploration as shown by Fortunato et al. (2017). Second, the function must be factorized to allow both the policy and the value function to condition the factors of the randomized reward function. This avoids the perception of the reward as partially observable by the policy. To fulfill these criteria, we implement the randomized reward function (denoted as $F$) as the dot product of the state feature and a randomly chosen latent vector:
$F(s,\boldsymbol{z})=\phi(s)\cdot\boldsymbol{z},$ (2)
where $\phi:\mathcal{S}\to\mathbb{R}^{d}$ is a feature extractor that transforms a state into a $d$-dimensional vector, and $\boldsymbol{z}\in\mathbb{R}^{d}$ represents a latent vector. Randomized rewards for each state are generated by sampling $\boldsymbol{z}$ from a given distribution $P_{\boldsymbol{z}}$, and then setting rewards as $F(s,\boldsymbol{z})$.
Latent conditioned policy and value network. Recall that the policy and the value function have to be aware of the state and the random variable that factorizes the randomized reward function $F$. To achieve this, we augment the input to the policy and the value functions with the latent vector $\boldsymbol{z}$. The resulting policy and the value networks are $\pi(.|s,\boldsymbol{z})$ and $V^{\pi}(s,\boldsymbol{z})$. We train the latent-conditioned value network to approximate the expected sum of the original reward and the randomized rewards as below
$V^{\pi}(s,\boldsymbol{z})\approx\mathbb{E}_{\pi}\left[\sum^{\infty}_{t=0}\gamma^{t}(R(s_{t},a_{t})+F(s_{t+1},\boldsymbol{z}))\right],$ (3)
and train the latent-conditioned policy $\pi$ to maximize $V^{\pi}(s,\boldsymbol{z})$ at every state $s$ and latent vector $\boldsymbol{z}$. Both value and policy networks can be trained with any off-the-shelf RL algorithms, e.g. PPO (Schulman et al., 2017), DQN (Mnih et al., 2015), A3C (Mnih et al., 2016), SAC (Haarnoja et al., 2018).
Latent vector sampling. To randomize the latent vector $\boldsymbol{z}$, rather than resampling it at each timestep, which could cause dithering actions due to fast-changing latent vectors fed to the policy, it is resampled at the start of each trajectory. This ensures each trajectory is rolled out under the same policy and latent vector $\boldsymbol{z}$, maintaining temporal consistency crucial for deep exploration, as indicated by prior studies (Osband et al., 2016a; Fortunato et al., 2017). The distribution of $\boldsymbol{z}$ will be elaborated in Section 4.
As we train the policy conditioned on the latent factor of the randomly sampled reward function, we term our method as Random Latent Exploration (RLE). We outline the algorithm in Algorithm 1 and present the detailed version in Algorithm 2. Note that at line 6 in Algorithm 1, we compute the randomized reward at the next state $s_{t+1}$ since the next state reflects the credit of taking action $a_{t}$ and state $s_{t}$. This choice is also common in prior works that compute exploration bonuses at each step (Burda et al., 2019).
1:  Input: Latent distribution $P_{\boldsymbol{z}}$
2:  repeat
3:     Sample a fresh latent vector: $\boldsymbol{z}\sim P_{\boldsymbol{z}}$
4:     for $t=0,\dots,T$ do
5:        Take action $a_{t}\sim\pi(.|s_{t},\boldsymbol{z})$ and transition to $s_{t+1}$
6:        Receive reward: $r_{t}=R(s_{t},a_{t})+F(s_{t+1},\boldsymbol{z})$
7:     end for
8:     Update policy network $\pi$ and value network $V^{\pi}$ with the collected trajectory $(\boldsymbol{z},s_{0},a_{0},r_{0},s_{1},\cdots,s_{T})$
9:  until convergence
Algorithm 1 Random Latent Exploration (RLE)
4 Experiments
We aim to show that our RLE improves over the action noise exploration method typically used in many RL algorithms (Schulman et al., 2017; Mnih et al., 2015) on the majority of tasks in both discrete and continuous control domains. Throughout the experiments, we train the agent for each task using PPO (Schulman et al., 2017) since it is a popular RL algorithm used in both discrete and continuous control tasks. Standard PPO implementation (Schulman et al., 2017) explores by sampling actions from the policy (i.e., a Boltzmann distribution over actions) learned.
We also compare RLE with the following exploration strategies as baselines:
  • NoisyNet (Fortunato et al., 2017): We chose it to be the representative baseline from the family of noise-based exploration (Osband et al., 2016a; Fortunato et al., 2017; Plappert et al., 2017) because it has been used in prior works on benchmarking exploration strategies (Chen et al., 2022; Taïga et al., 2019).
  • RND (Burda et al., 2019): We choose RND to be the representative baseline from the family of bonus-based exploration methods since it shows considerable improvements over action noises and noise-based approaches in hard-exploration tasks in Atari.
4.1 Illustrative Experiments on FourRoom
To test whether our method can perform deep exploration that is essential in many tasks requiring exploration, we start by running toy experiments on the the FourRoom environment (Sutton et al., 1999). We explain the setting of FourRoom environment and the results as follows.
Setup. Figure 1 illustrates FourRoom environment with $50\times 50$ states. The environment consists of four rooms separated by solid walls (which the agent can’t cross) and connected with small openings of a single state each (which the agent needs to go through to travel across rooms).
Figure 1: FourRoom environment. The agent starts at the top-right state (denoted by red ’S’) and can move left, right, up, and down. The black bars denote walls that block the agent’s movement.
The agent perceives the $(x,y)$ coordinates as the state input, and can take an action to move left, right, up, and down (if not interrupted by a wall). At the beginning of each trajectory, the agent always starts from the top-right corner of the room (denoted by the letter “S”). In this study, we always give zero rewards to the agent since we are interested in understanding and comparing how each exploration strategy behaves in the presence of sparse rewards, or even zero everywhere. This is also known as reward-free exploration.We also perform experiments on FourRoom with an sparse task reward of $1$ at the bottom-left corner. The results and visualization of visitation counts are deferred to Appendix C.
We compared the agents trained with different exploration strategies: PPO, NoisyNet, RND, and RLE. The corresponding policy in each of these exploration strategies is trained with Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017) for 2.5 million timesteps. For RLE, the feature extractor defined in Equation 2 is set to be a randomly initialized neural network with one hidden layer, the output layer of which has the same dimension as $\boldsymbol{z}\sim P_{\boldsymbol{z}}$. Further implementation details are available in Section B.2. We remark that, because of the walls, the FourRoom environment requires deep exploration to go to states distant from the initial states.
How does the latent vector influence the generated trajectories? The exploration of RLE is driven by sampling the latent vector $\boldsymbol{z}$ to change the behavior of the policy network $\pi$ and the objective of the value network $V^{\pi}$. One may wonder what is the impact of $\boldsymbol{z}$ on the induced policy’s behaviors. To evaluate this, we sample different latent vector $\boldsymbol{z}\sim P_{z}$ and rollout trajectories with the policy conditioning on those $\boldsymbol{z}$, plotting each trajectory in a different color in Figure 2. For this plot, we chose the checkpoint of the policy network stored in the middle of training (i.e., 1.5 million timesteps) since we want to study the behaviors of the policy network before it converges. Figure 2 demonstrates that the trajectories generated by different latent vectors $\boldsymbol{z}$ can span across all four rooms. This shows that altering the latent vector $\boldsymbol{z}$ can produce diverse trajectories and that latent vectors can influence the behaviors of the policy network. We provide the rollouts of trajectories for random seeds for RLE, PPO, RND and NoisyNet in Appendix C.
Figure 2: Rollout of multiple trajectories from a policy trained with RLE in the middle of the training (1.5 million timesteps), where each color denotes a distinct trajectory. As the figure demonstrates, changing the latent vector $\boldsymbol{z}$ in RLE leads to diverse trajectories across all four rooms.
Explaining the observed trajectory diversity. To explain why the trajectories generated by an RLE policy are diverse, we plot several of the reward functions induced by sampling different $\boldsymbol{z}$ in Figure 10. These plots demonstrate the diversity of the random rewards, each of which can guide the policy to a different part of the state space.
State Visitations in the environment. We plot the state visitation counts of each exploration strategies and present the results in Figure 3. The results shows that PPO’s state visitation centers around the initial room (i.e, top-right), indicating that action noises are not able to bring the agent far away from the initial state. In contrast, we see that RLE, RND, and NoisyNet are all able to frequently reach the rooms beyond the initial room, with RLE visitation count spread out across the four rooms. This suggests that RLE is capable of doing deep exploration similar to prior deep exploration algorithms for this environment.
Figure 3: State visitation counts of all the methods after training for 2.5M timesteps without any task reward (reward-free exploration). The start location is represented by the red ‘S’ at the top right. RLE achieves much wider state visitation coverage over the course of training compared to other baselines, confirming that the diverse trajectories generated by the policy are useful for exploration.
4.2 Benchmarking Results on Atari
Having performed illustrative experiments on the FourRoom toy environment, we now evaluate our method on more realistic and challenging tasks. We aim to show that RLE can improve PPO’s overall performance on most tasks.
Setup. We evaluate our method in the well-known Atari benchmark (Bellemare et al., 2013). Following the common practice in Atari (Mnih et al., 2015), the agent perceives a stack of the most recent four $84\times 84$ grayscale frames as inputs and takes discrete actions available in the gamepad of Atari (see Bellemare et al. (2013) for further environment details). For RLE, we chose the feature learned by the value network followed by a randomly initialized linear layer as $\phi$ (used in Equation 2) and set the dimension of the latent vector $\boldsymbol{z}$ as $8$. Note that the randomly initialized linear layer is frozen after initialization. We use the standard PPO hyperparameters provided in (Burda et al., 2019). The hyperparameters and implementation details of all the algorithms and PPO are deferred to Appendix B. For each Atari game (i.e., environment), we train each agent with $5$ different random seeds for $40$ million frames, as prior work (Chen et al., 2022; Bellemare et al., 2016) suggested. However, we trained Montezuma’s Revenge for $200$ million frames since its exploration difficulty is much harder than other Atari games (Burda et al., 2019).
Figure 4: Aggregated human normalized score across all 57 ATARI games. RLE exhibits a higher interquartile mean (IQM) of normalized score than PPO across 57 ATARI games, showing that RLE improves over PPO in the majority of tasks.
Does RLE improve the overall performance? We answer this question by calculating the interquartile mean (IQM) (Agarwal et al., 2021) and its 95% confidence interval, which was estimated using the bootstrapping method (DiCiccio & Efron, 1996) on the aggregated human-normalized scores from 57 games. Unlike empirical mean scores, IQM mitigates the influence of outliers on the aggregated metric. Figure 4 demonstrates that RLE achieves a higher IQM human-normalized score compared to all baselines, indicating that RLE enhances performance over other exploration strategies in the majority of Atari tasks. Besides the aggregate results, we present the learning curves for all methods across the 57 Atari games in Figure 18. Additionally, the final mean score of each method across five seeds for each Atari game is provided in Table 5 in the appendix. From Figure 18, we notice that RLE does not perform well on Montezuma’s Revenge, indicating that while RLE assists in producing diverse trajectories, Montezuma’s Revenge still presents a challenge when not relying on bonus-based exploration.
Does RLE improve over the baselines consistently? In addition to the margin of performance improvement on the aggregated score across all games, we demonstrate that RLE results in performance improvement over the baselines with high probability. Following the evaluation protocol suggested in Agarwal et al. (2021), we measure the probability of improvement (POI) between algorithms and their 95% confidence intervals, estimated using the bootstrapping method (DiCiccio & Efron, 1996), and present the results in Figure 5 and Figure 23. Figure 5(a) shows that the lower confidence bound of POI for RLE over each algorithm is above 0.5, indicating that RLE statistically outperforms the other baselines (Agarwal et al., 2021). This means that for a randomly chosen task in Atari, running RLE is likely to yield a higher score than the other baselines, implying that RLE’s performance improvements are consistent and not limited to a few games. Conversely, Figure 23(a) reveals that the POI over PPO for both NoisyNet and RND is below 0.5, suggesting that NoisyNet and RND do not consistently improve over PPO despite having better performance in a few games (see Figure 18). We use a CNN to represent the policy and value function in all Atari experiments, rather than an LSTM as done in (Chen et al., 2022) as the CNN-based architectures were also used in prior work (Burda et al., 2019) and due to its simplicity in implementation. This could explain why PPO outperforms RND over all 57 Atari games in our experiments.
Figure 5: (a) Probability of improvement (POI) of our method, RLE, over the baselines NoisyNet, RND and PPO across all $57$ Atari games (higher is better). The lower confidence bound of RLE’s POI over the other algorithms are all greater than $0.5$. This means that RLE statistically improves over other algorithms (Agarwal et al., 2021). (b) Probability of improvement of RLE over the baselines RND, PPO, and PPO with reward normalization across all $9$ IsaacGym tasks. In this domain as well, the lower confidence bound of RLE’s POI over the other algorithms are all greater than $0.5$. This means RLE statistically improves over the other algorithms.
4.3 Benchmarking Results on Isaac Gym
To demonstrate that RLE can improve upon PPO in both discrete and continuous control tasks, we also conducted experiments in IsaacGym (Makoviychuk et al., 2021), a benchmark suite containing numerous continuous control tasks. We compare the performance of RLE to that of PPO and RND as baselines. As RLE and RND use two reward streams with potentially different scales, we use reward normalization for these methods. Thus, we compare against PPO with reward normalization as an additional baseline.
We implemented RLE on top of PPO and trained it with standard PPO hyperparameters in IsaacGym provided in CleanRL, with implementation details provided in Section B.4. We selected both manipulation and locomotion tasks for evaluation and presented the learning curves in Figure 6, with learning curves on all environments presented in Figure 22.
Does RLE improve over baselines in continuous control? The results show that RLE achieves a higher average return than PPO in most tasks, with particularly large performance gains in AllegroHand and ShadowHand, indicating that RLE improves upon PPO in continuous control tasks. In AllegroHand and ShadowHand, the objective is to control a anthropomorphic hand to reorient objects to a target pose. These tasks require more exploration than other continuous control tasks since it takes many steps to achieve the target pose. Additionally, in Cartpole, PPO performance degrades abruptly in the middle of training, while RLE maintains high performance throughout, suggesting that RLE prevents the learning process from collapsing during training.
Does RLE achieve higher overall performance than PPO? To study the overall performance, we measure the probability of improvement of RLE over PPO and RND across 9 different IsaacGym tasks and also present the results in Figure 5 and Figure 23. Figure 5(b) shows that RLE has a statistically significant chance of improving over both PPO and RND in IsaacGym tasks as the lower confidence bound of the POI for RLE over each baseline method is greater than $0.5$. Furthermore, Figure 23(b) shows that out of all considered methods, RLE is the only one with a statistically significant POI over PPO.
We also measure the aggregated normalized return of RLE, PPO, and RND across all $9$ IsaacGym tasks. As the return is positive for PPO in each environment, we normalize runs by dividing by the mean score of PPO in that environment (i.e., normalize PPO to have a score of 1). For further details of this metric, see Section B.4.2. We present the results in Figure 24 and Figure 25, which show that RLE achieves higher aggregate performance than PPO and matches RND in this metric.
Figure 6: Comparison of performance between RLE, PPO, and RND in four IsaacGym tasks (higher is better). RLE achieves higher return than PPO and RND in the majority of tasks, especially in tasks like AllegroHand and ShadowHand that require more exploration. This suggests RLE improves over PPO and RND in continuous control domains as well.
4.4 Ablation Studies
We ran various ablation studies on Atari and IsaacGym environments to explore how different hyperparameters and design choices affect RLE performance.
Latent vector distribution. We investigated the impact of different latent vector distributions on RLE’s performance. Our study involves training RLE with various distributions, including $\mathrm{Uniform}([-0.5,0.5]^{d})$ and isotropic normal $\mathcal{N}(\boldsymbol{0},\mathbf{I}_{d})$ distributions, within a $d$-dimensional space where $d=8$. The detailed implementation is described in Appendix B. The results presented in Figure 7 indicate that RLE performs better than PPO across different latent vector distributions. This suggests that RLE’s efficacy is not significantly affected by the choice of latent vector distribution.
Figure 7: Performance of RLE with varying latent vector distribution $P_{\boldsymbol{z}}$ (see Section 3.1), where RLE (Sphere) is the one used in Section 4.2. The figure shows that RLE with the three distributions can all outperform PPO. This shows that RLE is not sensitive to the choice of latent vector distribution.
Latent vector dimension. This study explores how robust is RLE to different choices of the dimension $d$ of the latent vector z. We trained RLE for $d\in\{2,8,32,128\}$, where $d=8$ is the dimension used in the results presented in Section 4.2. The outcomes, depicted in Figure 8, demonstrate that RLE is capable of surpassing PPO across all tested dimensions. Although there are slight performance variations between different $d$ values, these differences are subtle, suggesting that RLE’s performance is insensitive to the choice of latent vector dimension $d$.
Figure 8: Performance of RLE with varying latent vector dimension $d$ (see Section 3.1), where RLE $(d=8)$ is the one used in Section 4.2. The figure shows that RLE can outperform PPO in all the four chosen dimensions. This shows that RLE is not sensitive to the choice of latent vector dimension.
Latent vector conditioning. In Section 3.1, we emphasized the necessity for the policy to be conditioned on the latent vector to prevent randomized rewards $F(s,\boldsymbol{z})$ from being perceived as noise by both the policy and the value network. This design choice’s importance is underscored by comparing RLE models with and without latent-conditioned policy and value networks, as shown in Figure 9. RLE without latent vector conditioning exhibited a performance drop in the Venture task, a hard-exploration game with sparse rewards. We hypothesize that the absence of latent vector conditioning results in limited behavioral variability in the policy network, as its outputs remain unchanged by different latent vector samples. This limitation likely leads to failures in challenging exploration tasks that necessitate a broader diversity in trajectory generation.
Figure 9: Performance of RLE with (w z-cond) and without (w/o z-cond) latent vector condition in the policy and value networks (see Section 3.1). The figure displays that RLE without latent vector condition suffers performance drop in Venture, a hard-exploration task with sparse rewards.
Feature vectors obtained from random neural network.
In the Atari experiments described in Section 4.2, we used a slow-moving estimate of the CNN features (see Section B.3.1) learned by the value network to compute RLE rewards. This choice of features slightly contributes to improved performance. Figure 20 presents the IQM of the normalized score and the POI of RLE over PPO with and without using the CNN features learned by the value network. The results demonstrate that incorporating value network features leads to a higher POI, while not significantly affecting the IQM. We plot the learning curve in each game for both variants in Figure 21, finding that while performance is broadly similar, there are a few games where the two variants have large differences in performance.
Choices of network architecture for the random reward network.
Since RLE relies on neural networks to extract features for computing random rewards $F(s,\boldsymbol{z})$, it’s important to examine how the choice of network architecture affects performance. We investigated the impact of different network architectures for extracting features $\phi(s)$ on the computation of RLE rewards $F(s,\boldsymbol{z})$ in IsaacGym (Makoviychuk et al., 2021), with the IQM of the normalized score in Figure 27 and POI over PPO shown in Figure 26. In our original IsaacGym experiments, we used the value network’s architecture for RLE. In this ablation study, we tested a shallower neural network architecture. The results indicate that RLE with a shallower network still performs well, suggesting that RLE is not highly sensitive to the choice of network architecture.
White noise random rewards.
In Section 3.1, we emphasized the importance of ensuring that RLE’s random reward function $F$ is correlated with states. If the random rewards are not state-dependent, they will act as white noise and will not enhance performance. In this ablation study, we empirically investigate the significance of making random rewards dependent on states. We compare the state-dependent RLE reward $F(s,\boldsymbol{z})$ with white noise rewards sampled from a normal distribution with zero mean and unit variance. This study was conducted in IsaacGym environments, with the results presented in Figure 28. The results demonstrate that white noise rewards significantly degrade performance, indicating that RLE rewards are not merely white noise.
5 Related Works
Random reward prediction was used as an auxiliary task Jaderberg et al. (2016) for improving representation learning in prior works (Dabney et al., 2021; Lyle et al., 2021). A closely related work is Ramesh et al. (2022), which employs a random general value function (GVF) (Sutton et al., 2011) for exploration by initializing a random reward function and using an ensemble of networks to predict policy-based random reward sums. The difference in prediction and the Monte Carlo estimate of random rewards, multiplied by prediction variance, enhances the agent’s reward. Our work presents a distinct approach from Ramesh et al. (2022) both in terms of motivation and implementation. Contrary to Ramesh et al. (2022), which aligns with previous studies (Burda et al., 2019; Pathak et al., 2017) by employing prediction errors as exploration bonuses, our RLE algorithm directly trains the policy with random rewards, demonstrating superior performance. This finding underscores that RLE provides a new angle to design exploration strategy beyond using prediction errors as exploration bonuses. Additionally, our RLE algorithm offers a more straightforward implementation by eliminating the need for ensemble training and Monte Carlo return estimation of random rewards. The detailed discussion on the relevant literature can be found in Appendix A.
6 Discussion and Conclusion
In this paper, we proposed a new exploration method called RLE that is straightforward to implement in practice, and effective in challenging deep RL benchmarks like Atari. We conclude with discussions and future work directions:
Simple plug-in for deep RL algorithms. RLE simply requires adding randomized rewards to the rewards received by the agent and augmenting the agent’s input with additional latent variables that correlate to these randomized rewards. As a result, RLE is agnostic to the base RL algorithm and can be integrated with any RL algorithm. Given its simplicity, generality, and the overall performance improvement it provides, we recommend using RLE as the default exploration strategy in deep RL implementations.
Connection to posterior sampling. At a high level, while RLE seems similar to the posterior sampling-based approaches (Thompson, 1933, 1935; Russo et al., 2018) in the sense that both utilize randomization for exploration, there are important differences: Firstly, the two methods explore via different mechanisms. Posterior sampling randomizes over different models of the environment, whereas RLE perturbs the reward function using random rewards. Secondly, the sampling distribution $P_{z}$ is fixed throughout learning in RLE, whereas the posterior distribution in posterior sampling changes with time and needs to be computed for every round (which is often challenging in practice). Thirdly, in posterior sampling, the posterior will eventually concentrate around the true model, and thus the algorithm will execute the optimal policy for the underlying environment. Whereas, in RLE, since the task rewards are constant throughout learning whereas random rewards change, in the later stage of the learning, the trained policy $\pi_{z}$ should focus on optimizing just the task rewards, and we believe that the random rewards will simply act as a regularization.
Benefits from parallelization. Note that, by design, our algorithm samples independent $\boldsymbol{z}$ in every round and can thus benefit from parallelization by running the algorithm on multiple copies of the same environment (when possible, e.g. using a simulator). Since different $\boldsymbol{z}$ produce diverse trajectories (see Figure 2 or Figure 14 for illustrations), multiple parallel copies of the same agent will simply produce more diverse data which would accelerate exploration.
On the inductive bias of $\boldsymbol{\phi}$. RLE is modular as one can choose any feature extractor $\phi(s)$ e.g. Transformer networks (Vaswani et al., 2017), MLPs, or even nonparametric models such as kernels. In our Atari experiments we use a CNN for $\phi$, but it would be interesting to explore how other choices of $\phi$ affect the diversity of the induced reward functions, and hence the generated trajectories.
$z$-sampling.
At every timestep in Algorithm 1 the latent variable $\boldsymbol{z}$ is sampled independently from the fixed distribution $P_{z}$ which is chosen at initialization. However, it is also intuitive to expect that $P_{z}$ should change as we learn more about the underlying environment. Looking forward it would be interesting to explore algorithms that change $P_{z}$ while training, e.g. to sample more from the set of latent variables which have not been explored yet or for which the corresponding policies $\pi_{z}$ have historically performed well in the given environment.
Other limitations. Currently, we limit our study to on-policy algorithms. Looking forward it would be interesting to extend RLE to off-policy algorithms such as DQN (Mnih et al., 2015) and SAC (Haarnoja et al., 2018); A practical way to do so would be to condition the $Q$-function on $\boldsymbol{z}$ in addition to its usual inputs. Separately, an important direction for future work is to explore the method in more continuous control and real-world robotics domains. While it is clear that our approach scales to high-dimensional state spaces in Atari and continuous control tasks in IsaacGym, it would also be interesting to see how it would generalize for real-world RL applications, e.g. in robotics. Finally, while RLE leads to performance improvements in general, we note that it does not improve performance in the famous hard-exploration game Montezuma’s Revenge. Thus, in the future, it would be interesting to extend RLE to environments and tasks with even sparser rewards.
Acknowledgement
We thank members of the Improbable AI Lab for helpful discussions and feedback. We are grateful to MIT Supercloud and the Lincoln Laboratory Supercomputing Center for providing HPC resources. This research was supported in part by Hyundai Motor Company, Quanta Computer Inc., MIT-IBM Watson AI Lab, an AWS MLRA research grant, ARO MURI under Grant Number W911NF-23-1-0277, ARO MURI under Grant Number W911NF-21-1-0328, and ONR MURI under Grant Number N00014-22-1-2740, NSF through award DMS-2031883, DOE through award DE-SC0022199, MIT UROP through the John Reed UROP Fund and the Paul E. Gray (1954) UROP Fund, and Simons Foundation. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the United States Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes, notwithstanding any copyright notation herein.
Author Contributions
  • Srinath Mahankali ran initial experiments to investigate the benefit of random rewards that informed the eventual formulation of RLE, which he then compared against baseline methods on Atari and IsaacGym environments and helped with paper writing.
  • Zhang-Wei Hong conceived the possibility of using random rewards for exploration. He was involved in research discussions, helped scale experiments, played a significant role in paper writing, and advised Srinath.
  • Ayush Sekhari was involved in research discussions and helped set the overall formulation of RLE. He played a significant role in paper writing, and advised Srinath and Zhang-Wei.
  • Alexander Rakhlin was involved in research discussions and advising.
  • Pulkit Agrawal was involved in research discussions, overall advising, paper writing, and positioning of the work.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. Our approach aims to accelerate RL in real-world domains, and depending on the domain to which it will be applied, we see many potential societal consequences of our work, however, none of them we feel must be specifically highlighted here.
Appendix A Additional Related Works
There are two main approaches in RL for exploration, (a) Randomize the agent by injecting noise thus inducing diverse behavior, and (b) Provide explicit reward bonuses that incentivize the agent to go to novel states. RLE, bridging the two approaches, injects noise into the agent by adding random latent reward bonuses during training.
Randomness is the key tool in many exploration strategies in RL. Perhaps the most popular example is $\epsilon$-greedy or Boltzmann sampling-based exploration (Mnih et al., 2013; Dann et al., 2022; Cesa-Bianchi et al., 2017; Eysenbach et al., 2018), which explores by playing random actions. Entropy regularization (Williams & Peng, 1991; Mnih et al., 2016), and MaxEnt RL (Haarnoja et al., 2018; Eysenbach & Levine, 2021; Garg et al., 2023; Hazan et al., 2019) are other instances of exploration algorithms that utilize randomness, as they explicitly bias towards learning policies that have a high entropy. Another exploration approach is to directly inject noise into the parameters of the policy or value networks, e.g. in off-policy methods  (Fortunato et al., 2017), RLHF with linear MDPs (Wu & Sun, 2023), online RL in tabular (Osband et al., 2016b), and linear MDPs (Zanette et al., 2020). Another line of work includes using Thompson sampling (Thompson, 1933, 1935; Russo et al., 2018), or posterior sampling for exploration (Osband et al., 2013; Gopalan & Mannor, 2015; Kveton et al., 2021; Zhang, 2022), which maintains a posterior distribution over the ground truth model and relies on the uncertainty in the posterior distribution for exploration. Posterior sampling, however, is intractable in practical RL settings due to the need to sample from the extremely complex posterior distribution; Various empirical approaches aim to sample from an approximate posterior instead (Li et al., 2021; Dwaracherla et al., 2020), but are unfortunately memory intensive. We note that RLE is different from these other works as it explores by adding random rewards instead of randomizing over the actions, policies, or models of the environment.
Exploration by adding explicit reward bonuses is also well studied in both theoretical and applied RL literature. A popular technique is to add novelty-based exploration bonuses that are constructed using prediction errors in transition dynamics  (Pathak et al., 2017, 2019; Ramesh et al., 2022) or outputs of randomly initialized target network RND (Burda et al., 2019), etc. Other approaches construct reward bonuses using upper confidence bounds on the uncertainty estimates for the underlying model (Auer et al., 2008; Vaswani et al., 2019), using discriminatively trained exemplar models to estimate novelty (Fu et al., 2017), or using elliptical potentials when the MDP has a linear parameterization (Jin et al., 2020; Zhang et al., 2022; Agarwal et al., 2020). Unfortunately, these methods often introduce additional components into the learning setups, e.g. additional neural networks for generating bonuses and the associated hyperparameters, which can make learning unstable. In contrast, RLE is much simpler to deploy as it adds random reward functions that are computed using features from the policy network.
RLE closely resembles the idea of Follow The Perturbed Leader (FTPL) developed in RL theory literature (Kveton et al., 2019b, a, 2020; Rakhlin & Sridharan, 2016; Dai et al., 2022). FTPL-based methods explore by adding carefully designed perturbations to the reward function that can guarantee optimism; since the perturbations are closely tied to the underlying modeling assumptions, FTPL-based methods are currently limited to restricted settings like linear bandits, tabular MDPs, etc. In contrast, RLE simply adds a random reward function sampled from a fixed distribution $P_{z}$, and is thus applicable in more practical RL settings. Another major difference is that our method also utilizes $z$-conditioned policies and value functions, and thus the randomness is shared amongst the reward function, policy network and value network.
Finally, there is a long line of work in the theoretical RL literature on developing exploration algorithms that can optimally exploit the structure of the underlying task in model-based RL (Ayoub et al., 2020; Sun et al., 2019; Foster et al., 2021), model-free RL (Jin et al., 2021; Du et al., 2021) and agnostic RL setting (Jia et al., 2023), however, the focus in these works is on statistical efficiency and the provided approaches are not computationally tractable.
Appendix B Experimental Implementation Details
In this section, we provide the hyperparameters and implementation details of our algorithm (RLE) along with the baseline methods (PPO, RND, NoisyNet) for the FourRoom and Atari Environments. We also provide hyperparameters and implementation details for all IsaacGym experiments.
B.1 RLE Pseudocode
Below, we provide the pseudocode for RLE in Algorithm 2.
1:  Input: Latent distribution $P_{\boldsymbol{z}},N$ parallel workers, $T$ steps per update, $S$ steps per sampling, feature network update rate $\tau$
2:  Randomly initialize a feature network $\phi$ with the same backbone architecture as the policy and value networks
3:  Initialize running mean $\boldsymbol{\mu}=\mathbf{0}$ and standard deviation $\boldsymbol{\sigma}=\mathbf{1}$ estimates of $\phi(s)$ over the state space
4:  Sample an initial latent vector for each parallel worker: $\boldsymbol{z}\sim P_{\boldsymbol{z}}$
5:  repeat
6:     Sample initial state $s_{0}$.
7:     for $t=0,\dots,T$ do
8:        Take action $a_{t}\sim\pi(.|s_{t},\boldsymbol{z})$ and transition to $s_{t+1}$
9:        Compute feature $f(s_{t+1})=(\phi(s_{t+1})-\boldsymbol{\mu})/\boldsymbol{\sigma}$
10:        Compute random reward: $F(s_{t+1},\boldsymbol{z})=\frac{f(s_{t+1})}{\|f(s_{t+1})\|}\cdot z$
11:        Receive reward: $r_{t}=R(s_{t},a_{t})+F(s_{t+1},\boldsymbol{z})$
12:        for $i=0,1,\dots,N-1$ do
13:           if worker $i$ terminated or $S$ timesteps passed without resampling then
14:              Resample sample $\boldsymbol{z}\sim P_{\boldsymbol{z}}$ for worker $i$
15:           end if
16:        end for
17:     end for
18:     Update policy network $\pi$ and value network $V^{\pi}$ with the collected trajectory $(\boldsymbol{z},s_{0},a_{0},r_{0},s_{1},\cdots,s_{T})$
19:     Update feature network $\phi$ using the value network’s parameters: $\phi\leftarrow\tau\cdot\pi+(1-\tau)\cdot\phi$
20:     Update $\boldsymbol{\mu}$ and $\boldsymbol{\sigma}$ using the batch of collected experience.
21:  until convergence
Algorithm 2 Detailed Pseudocode for Random Latent Exploration (RLE)
B.2 FourRoom Environment
We provide the hyperparameters used for experiments in the FourRoom environment in Table 1.
Parameter Value
PPO
      Total Timesteps $2,500,000$
      Optimizer Adam
      Learning Rate $0.001$
      Adam Epsilon $0.00001$
      Parallel Workers $32$
      Steps per Batch $128$
      Discount Rate $0.99$
      Generalized Advantage Estimation $\lambda$ $0.95$
      Minibatches per Epoch $4$
      Epochs per Training Step $4$
      Clipping Coefficient $0.2$
      Entropy Loss Weight $0.01$
      Discount Rate $0.99$
      Value Loss Weight $0.5$
      Gradient Norm Bound $0.5$
      Use Advantage Normalization True
      Use Clipped Value Loss True
      Policy Network Architecture MLP (64,64,4)
      Value Network Architectures MLP (64,64,1)
      Network Activation Tanh
NoisyNet
      Initial $\sigma$ $0.017$
RND
      Intrinsic Reward Coefficient $1.0$
      Drop Probability $0.25$
      Predictor Network Architecture MLP (256, 256, 256, 256, 256)
      Target Network Architecture MLP (64,256)
      Network Activation ReLU
RLE
      Intrinsic Reward Coefficient $0.1$
      Latent Vector Dimension $4$
      Feature Network Architecture MLP (64,64,64,4)
      Network Activation ReLU
Table 1: Hyperparameters and network architectures for FourRoom experiments.
B.2.1 RLE Implementation in FourRoom Enviroment
In our implementation of RLE for FourRoom environment, we ensure that the random reward functions $F(s,z)$ take values in $[-1,1]$.To compute the reward given a state $s$ and latent variable $\boldsymbol{z}$, we normalize the output of $\phi(s)$ to have unit norm. Specifically, we define the reward as:
$F(s,\boldsymbol{z})=\frac{\phi(s)}{\|\phi(s)\|}\cdot z,$
where $\phi$ is the randomly initialized feature network that transforms the state $s$ to a vector with the same dimension as $\boldsymbol{z}$. In the FourRoom environment, we sample $\boldsymbol{z}$ from the unit sphere at every training step, which occurs every $128$ timesteps. We perform the sampling independently for each of the $32$ parallel workers.
B.3 Atari
We display the hyperparameters used for experiments in Atari games in Table 2. For PPO and RND, we use the default hyperparameters based on the cleanrl codebase (Huang et al., 2022), which were tuned for Atari games. For NoisyNet, we use the same hyperparameters as PPO with the exception of the entropy loss weight, which is set to 00 as recommended by (Fortunato et al., 2017). We give a detailed description of the Atari implementation of RLE below.
Parameter Value
PPO
      Total Timesteps $40,000,000$
      Optimizer Adam
      Learning Rate $0.0001$
      Adam Epsilon $0.00001$
      Parallel Workers $128$
      Steps per Batch $128$
      Discount Rate $0.99$
      Generalized Advantage Estimation $\lambda$ $0.95$
      Minibatches per Epoch $4$
      Epochs per Training Step $4$
      Clipping Coefficient $0.1$
      Entropy Loss Weight $0.01$
      Discount Rate $0.99$
      Value Loss Weight $0.5$
      Gradient Norm Bound $0.5$
      Use Advantage Normalization True
      Use Clipped Value Loss True
      Policy Network Architecture CNN + MLP (256,448,448,18)
      Value Network Architectures CNN + MLP (256,448,448,1)
      Network Activation ReLU
NoisyNet
      Initial $\sigma$ $0.017$
      Entropy Loss Weight 00
RND
      Intrinsic Reward Coefficient $1.0$
      Extrinsic Reward Coefficient $2.0$
      Update Proportion $0.25$
      Observation Normalization Iterations $50$
      Discount Rate $0.999$
      Entropy Loss Weight $0.001$
      Intrinsic Discount Rate $0.99$
      Predictor Network Architecture CNN + MLP (512,512,512)
      Target Network Architecture CNN + MLP (512)
      Network Activation LeakyReLU
RLE
      Intrinsic Reward Coefficient $0.01$
      Latent Vector Dimension $16$
      Latent Vector Resample Frequency $1280$
      Learning Rate $0.0003$
      Discount Rate $0.999$
      Intrinsic Discount Rate $0.99$
      Feature Network Update Rate $\tau$ $0.005$
      Feature Network Architecture CNN + MLP (256,448, 16)
      Network Activation ReLU
Table 2: Hyperparameters and network architectures for Atari experiments.
B.3.1 RLE Implementation Details in Atari
Feature network architecture and update.
We start with a randomly initialized neural network $\phi$ which takes a state $s$ as input and outputs a vector in $\mathbb{R}^{d}$, which has the same dimension as $\boldsymbol{z}$. In our implementation, $\phi$ contains a CNN backbone with an identical architecture to the (shared) policy and value backbone, along with a final linear layer on top to convert it to a low dimension $\mathbb{R}^{d}$. In our implementation, we choose $d=8$. To update $\phi$, we follow the rule:
$\mathrm{CNN}_{\phi}\leftarrow\tau\cdot\mathrm{CNN}_{V}+(1-\tau)\cdot\mathrm{CNN}_{\phi}$
for a small value of $\tau$, and we choose $\tau=0.005$ for our experiments. This network update is inspired by the target network update in DQN (Mnih et al., 2015) and does not require any gradient steps.
Computation of random reward bonus.
When the agent experiences a transition $(s,a,s^{\prime})$, we obtain random reward bonus from $\phi$ as follows: Obtain the low-dimensional vector output ϕ(s)italic-$\phi(s^{\prime})$. We standardize the output of ϕ(s)italic-$\phi(s^{\prime})$ using a running mean and standard deviation estimate so that the output is a normal distribution on $\mathbb{R}^{d}$. Meanwhile, sample a vector $\boldsymbol{z}\sim\mathcal{S}^{d-1}$, and compute the following value
$F(s^{\prime},\boldsymbol{z})=\frac{\phi(s^{\prime})\cdot\boldsymbol{z}}{\|\phi(s^{\prime})\|}.$
Policy input.
The policy observes the observation returned by the environment, which is $4$ stacked grayscale frames of dimension $(84,84)$. In addition, the policy observes $\boldsymbol{z}$ as well as the random reward $F(s_{t},\boldsymbol{z})$ from the previous time step.
Resampling of the latent variable $\boldsymbol{z}$.
In our Atari experiments, there are 128 parallel workers. We sample $\boldsymbol{z}$ independently across all workers from the $d$-dimensional unit sphere $\mathcal{S}^{d-1}$, and resample upon either of the following signals:
  1. 1.
    An environment has reached the ‘DONE’ flag, or
  2. 2.
    An environment has survived with this $\boldsymbol{z}$ for $1280$ time steps.
Policy training.
We use PPO with the augmented observation space train on the combined reward as usual. As we resample $\boldsymbol{z}$ during an episode, we also treat the problem of maximizing randomized reward as episodic. Specifically, we set the ‘done’ signal to True whenever we resample $\boldsymbol{z}$. Thus, we do not use returns from future $\boldsymbol{z}$ within the same episode to estimate the return under the current $\boldsymbol{z}$.
B.3.2 Evaluation Details in Atari
Human normalized score
To compute aggregate performance, we first compute the human normalized score for each seed in each environment as $\frac{\mathrm{Agent}_{\mathrm{score}}-\mathrm{Random}_{\mathrm{score}}}{\mathrm{Human}_{\mathrm{score}}-\mathrm{Random}_{\mathrm{score}}}$. After this, we compute the IQM to measure aggregate performance as recommended in (Agarwal et al., 2021) as it is robust to outliers.
Capped human normalized score
We use the capped human normalized score (CHNS) (Badia et al., 2020) to measure the aggregate performance of RLE and baselines in Figure 19. To compute the CHNS, we first compute the human normalized score (HNS) of the agent, as done in (Badia et al., 2020), as $\frac{\mathrm{Agent}_{\mathrm{score}}-\mathrm{Random}_{\mathrm{score}}}{\mathrm{Human}_{\mathrm{score}}-\mathrm{Random}_{\mathrm{score}}}$, after which it is clipped to be between 00 and $1$. In addition to aggregate metrics, we provide individual mean scores of all methods in all 57 games in Table 5 along with the corresponding learning curves in Figure 18.
Probability of improvement
We use the probability of improvement (POI), recommended in (Agarwal et al., 2021), to measure the relative performance between algorithms across all 57 Atari games.
Bootstrapped confidence intervals
We use the bootstrapping method (DiCiccio & Efron, 1996; Agarwal et al., 2021) to estimate the confidence intervals for all aggregated metrics we report, and mean performance for an algorithm in one environment.
B.4 IsaacGym
We display the hyperparameters used for experiments in IsaacGym in Table 3. For PPO, we use the default hyperparameters recommended by the cleanrl codebase (Huang et al., 2022), which were tuned for IsaacGym tasks and can vary across tasks (specifically, using different hyperparameters for the ShadowHand and AllegroHand tasks). For RLE, we use the same hyperparameters for each task. We also display the environment-specific hyperparameters in Table 4, which are shared for each training algorithm we consider in our experiments.
Parameter Value
PPO
      Optimizer Adam
      Learning Rate $0.0026$
      Adam Epsilon $0.00001$
      Steps per Batch $16$
      Discount Rate $0.99$
      Generalized Advantage Estimation $\lambda$ $0.95$
      Minibatches per Epoch $2$
      Epochs per Training Step $4$
      Clipping Coefficient $0.2$
      Entropy Loss Weight $0.0$
      Discount Rate $0.99$
      Value Loss Weight $2.0$
      Gradient Norm Bound $1.0$
      Use Advantage Normalization True
      Use Clipped Value Loss False
      Policy Network Architecture MLP (256,256,256)
      Value Network Architecture MLP (256,256,256,1)
      Network Activation Tanh
      Reward Scale $1.0$
RND
      Intrinsic Value Loss Weight $2.0$
      Intrinsic Reward Coefficient $0.5$
      Update Proportion $0.0625$
      Observation Normalization Iterations $50$
      Predictor Network Architecture MLP (256, 256, 256, 256)
      Target Network Architecture MLP (64, 64, 256)
      Network Activation ReLU
RLE
      Intrinsic Value Loss Weight $0.5$
      Intrinsic Reward Coefficient $0.01$
      Latent Vector Dimension $32$
      Latent Vector Resample Frequency $16$
      Learning Rate $0.001$
      Feature Network Update Rate $\tau$ $0.005$
      Policy Network Architecture MLP (256,256,256)
      Value Network Architecture MLP (512,512,256,256,1)
      Feature Network Architecture MLP (512,512,256)
      Network Activation Tanh
PPO (AllegroHand and ShadowHand)
      Steps per Batch $8$
      Minibatches per Epoch $4$
      Epochs per Training Step $5$
      Reward Scale $0.01$
Table 3: Hyperparameters and network architectures for IsaacGym experiments. The number of training steps and parallel workers depends on the environment, but are shared across different methods.
Parameter Value
AllegroHand
      Number of Timesteps $600,000,000$
      Number of Parallel Environments $8,192$
ShadowHand
      Number of Timesteps $600,000,000$
      Number of Parallel Environments $8,192$
BallBalance
      Number of Timesteps $200,000,000$
      Number of Parallel Environments $4,096$
Humanoid
      Number of Timesteps $200,000,000$
      Number of Parallel Environments $4,096$
Ant
      Number of Timesteps $100,000,000$
      Number of Parallel Environments $4,096$
Cartpole
      Number of Timesteps $100,000,000$
      Number of Parallel Environments $4,096$
FrankaCabinet
      Number of Timesteps $100,000,000$
      Number of Parallel Environments $4,096$
Anymal
      Number of Timesteps $100,000,000$
      Number of Parallel Environments $4,096$
AnymalTerrain
      Number of Timesteps $100,000,000$
      Number of Parallel Environments $4,096$
Table 4: Environment-specific parameters and their values. These parameters are shared across all algorithms.
B.4.1 RLE Implementation Details in IsaacGym
Feature network architecture and update
Similar to our implementation of RLE in the Atari domain, we start with a randomly initialized neural network $\phi$ that has an MLP backbone with the same architecture as the backbone of the value function. We update the backbone paramaters using the same slow moving average as in Section B.3.1 with $\tau=0.005$:
$\mathrm{MLP}_{\phi}\leftarrow\tau\cdot\mathrm{MLP}_{V}+(1-\tau)\cdot\mathrm{MLP}_{\phi}.$
Computation of random reward bonus
We standardize the output of ϕ(s)italic-$\phi(s^{\prime})$ using a running mean and standard deviation estimate so the output approximates a normal distribution on $\mathbb{R}^{d}$. We sample a vector $\boldsymbol{z}\sim\mathcal{S}^{d-1}$ and compute the reward as:
$F(s^{\prime},\boldsymbol{z})=\phi(s^{\prime})\cdot\boldsymbol{z}.$
Note that this is slightly different from the implementation in Atari, where we divide by ϕ(s)normitalic-$\|\phi(s^{\prime})\|$. We use reward normalization for RLE in both domains to scale the randomized reward, so both types have a similar effect.
B.4.2 Evaluation Details in IsaacGym
PPO normalized score
We use the IQM of the PPO normalized score to compute aggregate performance across 9 different environments in IsaacGym. We compute the PPO normalized score of the agent as $\mathrm{Agent}_{\mathrm{score}}/\mathrm{PPO}_{\mathrm{mean}}$. For example, the mean performance of PPO in a single environment under the PPO normalized score will be $1$. We compute the IQM of this metric for 5 seeds across 9 games (or 45 total runs) to aggregate performance.
Other evaluation details
Similar to our experiments in the Atari domain, we use the bootstrapping method to estimate confidence intervals and use the probability of improvement to measure relative performance between different algorithms.
Appendix C Visualizations on FourRoom
In this section, we provide further results and visualizations for the FourRoom environment:
Figure 10: Visualization of the reward function $F(s;\boldsymbol{z})$ for 10 different random choices of $\boldsymbol{z}$ in FourRoom environment. The reward is is given by $F(s,\boldsymbol{z})=\boldsymbol{z}\cdot\phi(s)/\|\phi(s)\|$. The above image demonstrates the diversity and coverage of random reward functions in the FourRoom environment.
(a) After 500K timesteps.
(b) After 2.5M timesteps.
Figure 11: State visitation counts for different methods on FourRooms environment trained for 500K and 2.5M timesteps with task reward. The start location is the top-right state of the grid (represented by the red ‘S’). The agent gets a task reward of $1$ at the bottom-left state (represented by red ‘*’).
RL with task reward.
In addition to the reward-free setting, we train all methods in the FourRoom environment in a sparse-reward setting for 2.5M timesteps. There is a reward of $1$ in the bottom-left corner, and the reward is 00 at all other states. We plot the state visitation counts of all methods after 500K and 2.5M timesteps in Figure 11. In addition, we train five seeds in this environment for each method, and find that the average score for RLE and NoisyNet is $0.6$, while the average score for RND and PPO is 00. This suggests that the FourRoom environment is a task that requires exploration as it is difficult for methods that rely on action noise like PPO to achieve any reward.
State visitations.
  • Figure 11 shows state visitation counts for all algorithms trained with a sparse task reward which is $1$ at the bottom-left state (red ’*’) and 00 everywhere else.
  • Figure 12 shows state visitation counts for all algorithms trained for 500K and 2.5M steps without any task reward.
  • Figure 13 shows state visitation counts for RLE trained in a modified version of the environment with stochastic observations within a 2x2 square region of the environment. Through this, we test if RLE is susceptible to the “NoisyTV” problem (Burda et al., 2019).
Visualizations of trajectory diversity across algorithms.
  • Figure 14 shows 5 trajectories sampled from policies trained with RLE across 5 different seeds at three different points in training: after 500K steps, 1.5M steps, and 2.5M steps.
  • Figure 15 shows the same for NoisyNet.
  • Figure 16 shows the same for RND.
  • Figure 17 shows the same for PPO.
From visual evaluation, the above plots suggest that RLE induces more diverse trajectories as compared to other baselines (PPO, RND, and NoisyNet) on the FourRoom environment.
(a) After 500K steps
(b) After 2.5M steps
Figure 12: State visitation counts for different algorithms on FourRooms environment after training for 500K timesteps and 2.5M timesteps. All algorithms were trained without task reward (reward-free exploration).
Figure 13: State visitation counts for RLE when trained in an environment with stochasticity in the observation space. The observation is only stochastic within the red square and is deterministic everywhere else. Even after discovering the red square, the agent is able to discover states outside of those regions and continues to explore throughout training. This suggests that RLE is less affected by the NoisyTV problem compared to novelty-based exploration methods.
Figure 14: Visualization of trajectories generated by sampling from a policy trained with RLE for 2.5M timesteps in a reward-free setting across 5 seeds at different points in training. We sample 5 trajectories for each seed.
Figure 15: Visualization of trajectories generated by sampling from a policy trained with NoisyNet for 2.5M timesteps in a reward-free setting across 5 seeds at different points in training. We sample 5 trajectories for each seed.
Figure 16: Visualization of trajectories generated by sampling from a policy trained with RND for 2.5M timesteps in a reward-free setting across 5 seeds at different points in training. We sample 5 trajectories for each seed.
Figure 17: Visualization of trajectories generated by sampling from a policy trained with PPO for 2.5M timesteps in a reward-free setting across 5 seeds at different points in training. We sample 5 trajectories for each seed.
Appendix D Detailed Results and Learning Curves on all Atari Games
We provide:
  • The scores for each of the algorithms (RLE (Ours), PPO, RND and NoisyNet) on all 57 Atari games in Table 5. Each of the algorithms was trained for 40M steps on all Atari games except for the results for Montezuma’s Revenge where we trained for 400M steps. The reported Human performance is obtained from Mnih et al. (2013); Badia et al. (2020).
  • Learning curves for all the algorithms (RLE (Ours), PPO, RND and NoisyNet) for all 57 Atari games in Figure 18.
  • Aggregated capped human normalized score (described in Section B.3.2) for each of the algorithms (RLE (Ours), PPO, RND and NoisyNet) over all 57 Atari games.
  • An ablation study of how the soft update rule for the feature network affects performance on Atari games taking three metrics into account (scores on individual games, IQM of human normalized score, and probability of improvement over PPO).
Figure 18: Game scores for different algorithms for all 57 Atari games.
PPO RND NoisyNet RLE (Ours)
Alien-v5 1409.22 1010.94 1112.59 1680.13
Amidar-v5 595.53 304.01 210.88 836.01
Assault-v5 1886.68 7045.97 1605.63 8368.57
Asterix-v5 3392.84 2439.56 2892.50 5350.79
Asteroids-v5 2913.15 3232.00 3513.02 1798.64
Atlantis-v5 915703.86 957537.41 972286.70 979023.58
BankHeist-v5 1017.55 250.91 678.44 1036.61
BattleZone-v5 35832.79 32536.65 30021.64 34793.65
BeamRider-v5 5200.10 6915.42 5755.02 5895.93
Berzerk-v5 845.95 931.95 812.45 1445.49
Bowling-v5 46.84 67.17 56.17 35.88
Boxing-v5 97.99 67.44 97.06 93.94
Breakout-v5 227.34 139.41 127.66 337.07
Centipede-v5 4634.04 7972.15 4419.47 4151.10
ChopperCommand-v5 4342.80 2140.08 4787.63 9710.00
CrazyClimber-v5 112278.70 107602.05 116637.17 115593.01
Defender-v5 46957.65 43387.18 34448.07 47872.91
DemonAttack-v5 30714.39 44342.27 28287.09 45217.82
DoubleDunk-v5 -1.57 -5.04 -1.43 -1.51
Enduro-v5 387.89 595.65 150.38 990.95
FishingDerby-v5 31.13 -57.88 14.14 30.68
Freeway-v5 25.83 21.19 32.40 32.49
Frostbite-v5 949.08 2944.37 1747.85 4658.90
Gopher-v5 1020.40 11822.26 1055.82 13290.12
Gravitar-v5 920.19 597.42 674.99 1381.69
Hero-v5 25495.80 14695.30 11433.06 9668.68
IceHockey-v5 -2.09 -16.70 -1.34 -2.39
Jamesbond-v5 3157.81 9347.30 4633.37 2452.21
Kangaroo-v5 6504.67 5474.45 1596.99 6992.13
Krull-v5 8731.23 7264.60 9063.52 8981.43
KungFuMaster-v5 26131.84 30902.44 43341.34 27813.32
MontezumaRevenge-v5 2077.03 4406.79 0.00 79.48
MsPacman-v5 2417.82 1446.32 2127.62 2676.20
NameThisGame-v5 9392.45 6078.34 7818.62 13701.36
Phoenix-v5 7137.14 19195.54 4786.92 11272.80
Pitfall-v5 -0.67 -3.41 -0.05 -57.65
Pong-v5 14.52 -10.34 7.10 17.17
PrivateEye-v5 98.34 87.24 95.55 97.79
Qbert-v5 12168.36 4300.73 3381.40 16261.59
Riverraid-v5 9268.85 4267.51 5642.73 12009.63
RoadRunner-v5 30354.36 19452.68 27037.68 53920.12
Robotank-v5 28.68 22.11 26.34 36.71
Seaquest-v5 1172.22 2463.42 920.71 1724.96
Skiing-v5 -16370.14 -10644.07 -16398.72 -13887.77
Solaris-v5 2203.41 1206.94 2584.66 2203.76
SpaceInvaders-v5 938.00 878.91 981.24 1981.37
StarGunner-v5 52219.39 23174.16 42645.43 64011.13
Surround-v5 -3.41 -7.28 -2.41 -3.91
Tennis-v5 -2.03 -19.44 -1.06 -4.49
TimePilot-v5 8319.51 9695.60 10888.94 11636.24
Tutankham-v5 204.57 140.97 142.70 209.23
UpNDown-v5 212171.29 251442.66 219951.36 151036.48
Venture-v5 401.20 969.00 0.06 782.98
VideoPinball-v5 32654.24 35275.58 28236.75 84825.64
WizardOfWor-v5 8355.05 10151.69 6306.86 9942.29
YarsRevenge-v5 74833.17 71789.37 65902.61 58507.98
Zaxxon-v5 17354.21 6273.86 6104.86 17403.15
Table 5: Performance on all 57 Atari games. Each algorithm was trained for 40M timesteps, except for Montezuma’s Revenge where we trained for 400M timesteps. The reported Human performance is obtained from Mnih et al. (2013); Badia et al. (2020).
Figure 19: Capped human normalized score across all 57 Atari games. RLE outperforms all other methods in this metric and requires half the training time to reach the same score as the next best method (PPO).
(a) Probability of improvement over PPO with and without a slow value feature update rule. Using the value features leads to a slight increase in performance.
(b) IQM of human normalized score of RLE, both with and without a slow value feature update rule. With respect to this metric, both versions of the method perform very similarly overall.
Figure 20: Comparison of RLE performance with and without a slow value feature update rule.
Figure 21: Game scores for our method with and without a slow value feature update rule. Performance is usually similar, but noticeably different in a handful of games (for example Alien, Frostbite, Skiing, Tutankham, YarsRevenge).
Appendix E Detailed Results and Learning Curves on all IsaacGym Tasks
We provide:
  • The learning curves for PPO and RLE in all 9 IsaacGym tasks that we consider in Figure 22.
  • The POI of all methods over PPO over all 9 IsaacGym tasks in Figure 23.
  • The IQM of PPO normalized score of RLE, PPO, and RND aggregated across all 9 IsaacGym tasks in Figure 24.
  • The mean of PPO normalized score of RLE, PPO, and RND aggregated across all 9 IsaacGym tasks in Figure 25.
  • An ablation study of how different network architectures for $\phi$ affect performance on IsaacGym tasks IQM of PPO normalized score, and probability of improvement over PPO). We plot the IQM of PPO normalized score in Figure 27, and probability of improvement over PPO in Figure 26.
  • An ablation study of how using white noise for randomizing rewards affects performance, shown in Figure 28.
Figure 22: Comparison of achieved returns between RLE and standard PPO (higher is better). RLE achieves return greater than or equal to that of standard PPO in the majority of tasks. We also compare RLE to an ablation of PPO that uses reward normalization and find that RLE improves over it as well. Finally, we compare RLE to RND, finding that while RND surprisingly improves performance in these tasks compared to PPO, RLE improves over RND in the majority of tasks.
Figure 23: (a) Probability of improvement of RLE, RND, and NoisyNet over PPO across all $57$ Atari games. POI over PPO of both NoisyNet and RND are below $0.5$, implying that neither NoisyNet nor RND statistically improve over PPO overall across $57$ Atari games. (b) Probability of improvement of RLE, RND, and PPO with reward normalization over PPO across all $9$ IsaacGym tasks. POI over PPO of both RND and PPO with reward normalization are below $0.5$, demonstrating that neither baseline statistically improves over PPO overall across the IsaacGym tasks.
Figure 24: Normalized score across all 9 IsaacGym tasks that we consider, aggregated using the IQM. RLE achieves a higher interquartile mean of normalized score compared to both variants of PPO, indicating that it can improve over PPO in continuous control domains as well. Meanwhile, RLE performs similarly to RND in terms of aggregated performance.
Figure 25: Mean normalized score across all 9 IsaacGym tasks. Aggregating using the mean yields similar results: RLE outperforms both variants of PPO, and slightly outperforms RND overall.
Figure 26: Probability of improvement over PPO of RLE and RLE with a different architecture for $\phi$. The probability of improvement for both RLE variants is close and the confidence intervals for the probability of improvement metric heavily overlap. This suggests that RLE is robust to the choice of architecture for $\phi$.
Figure 27: IQM of PPO normalized score of RLE and RLE with a different architecture for $\phi$. The different architecture used in this experiment has less width and uses one less layer. The IQM of normalized score is similar for both methods, suggesting that RLE does not highly depend on the architecture of the network $\phi$.
Figure 28: Comparison of achieved returns between RLE and PPO with random normal noise sampled i.i.d from a standard normal distribution added to the reward at each timestep. The intrinsic reward coefficient is $0.01$. RLE outperforms this variant of PPO in a large majority of games, suggesting that RLE benefits from using state-dependent random rewards.